OpenStudy (anonymous):

derivative of sqrt1+2x-sqrt1-4x/x

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

\[\frac{ \sqrt{1+2x}-\sqrt{1-4x} }{ x }\]

OpenStudy (anonymous):

use quotient rule

OpenStudy (anonymous):

I did but I still didn't get the right answer which is \[\frac{ 1 }{ \sqrt{1+2x}} +\frac{ 2 }{ \sqrt{1-4x}}\]

OpenStudy (anonymous):

got it thanx for link

OpenStudy (anonymous):

welcome